OpenStudy (anonymous):

Write the equation of the line with a slope of 3/7 passing through the points (9,-2)

OpenStudy (lgbasallote):

use point slope form \[y - y_1 = m(x-x_1)\] where m=slope does that help? or do you need more hints?

OpenStudy (anonymous):

the fraction confuses me i get how to set it up tho

OpenStudy (lgbasallote):

hmm okay set it up...then i'll tell you the next step

OpenStudy (anonymous):

y-(-2)=3/7(x-9)

OpenStudy (lgbasallote):

so simplifying that's \[y + 2 = \frac 37 (x-9)\] right?

OpenStudy (anonymous):

then i get lost after that.

OpenStudy (lgbasallote):

ahh you need to get rid of the denominator...you do this by multiplying EACH term by the denominator..in this case 7 so multiply both sides by 7

OpenStudy (anonymous):

7y+14=3x-27 right?

OpenStudy (lgbasallote):

correct

OpenStudy (lgbasallote):

now turn it into general form \[ax + by + c = 0\]

OpenStudy (anonymous):

would the answer be 3y-7y=41 ?

OpenStudy (lgbasallote):

dont you mean 3x - 7y - 41 = 0

OpenStudy (anonymous):

yes but i just move the -42 over

OpenStudy (anonymous):

41*

OpenStudy (lgbasallote):

well if in the choices it says = 41 then choose that...but in standard form it';s what i wrote

OpenStudy (anonymous):

ok thanks alot!! :)